NH3 is a weak base (Kb = 1.8 × 10–5) and so the salt NH4Cl acts as a weak acid. What is the pH of a solution that is 0.048 M in NH4Cl at 25 °C?

pH=_______?
Answer #1

The following reaction takes place in the solution :

               NH4+ + H2O ----> NH3 + H3O+

Initial         0.048                   0         0

Eqb.         0.048-x                x          x

Hydrolysis constant, Kh = [NH3] [H3O+] / [NH4+]

Also, Kh = Kw / Kb

Putting values we get :

Kh = 10-14 / 1.8*10-5 = 5.5*10-9

Putting this in above equation we get :

x2/0.048-x = Kh

Since Kh is small and x is also supposed to be small, we can safely ignore the term x*Kh.

Solving further we get :

x = 5.164*10-6

pH = -log10[H3O+] = -log x = 5.29